OpenStudy (anonymous):

how do i make y = 2-(4/(X+1)) be in the form x= ...

OpenStudy (anonymous):

Start moving the things on the right of the = to the left one at a time, example. y = 2-(4/(X+1)) y + 2 = 2-(4/(X+1)) +2 y + 2 = -(4/(X+1)) Does that make sense?

OpenStudy (anonymous):

yes but thats where im stuck

OpenStudy (anonymous):

OK, well if you were going to take a stab at it, what would you try next?

OpenStudy (anonymous):

could you show me hoe to work it out from there

OpenStudy (anonymous):

Next try ignoring the left side for a moment. How would you get rid of the "-" sign in -(4/(x+1)) Thin of it as -1*(4/(x+1)

OpenStudy (anonymous):

Thin = Think, sorry.

OpenStudy (unklerhaukus):

y = 2-(4/(x+1)) taking away 2 from both sides of the equation y - 2 = 2-(4/(x+1)) -2 simplifying the right hand side y - 2 = -(4/(x+1))

OpenStudy (unklerhaukus):

now you can multiply (or divide) both sides of the equation by -1